Frank has half as many again as sally who has half again as many again as mary. Altogether they have 209.
how many has each?

Answer by stanbon(75887)

Frank has half as many again as sally who has half again as many again as mary. Altogether they have 209.
how many has each?
---
Let Mary have "x"
Then Sally has "(3/2)x"
And Frank has (3/2)(3/2 x) = (9/4)x
-----------------------------------------
Equation:
x + (3/2)x + (9/4)x = 209
4x + 6x + 9x = 4*209
19x = 4*209
x = 44
(3/2)x = 66
(9/4)x = 99
==================
cheers,
Stan H.
==============